# Efficient Tensor Completion Algorithms for Highly Oscillatory Operators

### v2

@@ 3.3 Butterfly Decomposition in Machine Learning
Due to its ability to approximate Fourier integral operators with a small number of parameters, butterfly decomposition has recently been integrated into deep neural networks to efficiently learn structured linear transforms [16, 48]. Building on this, the Kaleidoscope hierarchy, which is a product of learnable orthogonal butterfly matrices, was introduced into the Mel-Frequency Spectral Coefficients (MFSC) pipeline [18], achieving performance comparable to hand-crafted transforms. These Kaleidoscope matrices are capable of learning latent permutations and accelerating inference [17]. To further enhance their practicality and compatibility with GPU architectures, the sparsity structure of the butterfly factors was modified in various works [52, 15, 10]. Empirical evaluations in these works show that restructuring maintains similar accuracy while improving efficiency by enhancing the ease of implementation.
Integrating butterfly matrices into machine learning frameworks like PyTorch has led to new developments in neural network architectures and computational kernels. Although butterfly decompositions theoretically enable matrix-vector products in $O(n\log n)$ complexity, achieving these speedups in practice poses significant challenges. Traditional block-sparse representations often lead to inefficient high-level implementations, as they require manual looping over sparse blocks, which impedes performance. Realizing theoretical speedups for both the forward pass, that is, applying the butterfly matrix, and backward pass, which is obtaining the gradients of block sparse matrices, demands carefully optimized software kernels. While recent efforts have developed efficient GPU kernels for the forward pass [28, 24], kernels for computing gradients (i.e., the backward pass) remain underdeveloped. To address this limitation, some works have modified the sparsity structure of the butterfly factors to better align with hardware efficiency [15]. In contrast, we propose a new representation of the butterfly decomposition as a tensor network, enabling differentiation with respect to the factors by leveraging automatic differentiation with tensor contractions. This reformulation is detailed in the next section.
@@ 4.1 Butterfly Decomposition Rewritten as a Tensor Network
To describe the tensor decomposition reformulation, we will continue to use the same assumptions as in the previous section. We consider an $m\times n$ input matrix $\bm{T}$, where $n=c\cdot 2^{L}$, and $L$ is the number of levels in the butterfly decomposition. The row and column sets, $\chi=\{0,\dots,m-1\}$ for rows and $\xi=\{0,\dots,n-1\}$ for columns of the matrix, and $m=n$. These sets are organized into hierarchical trees $\mathcal{T}_{\chi}$ and $\mathcal{T}_{\xi}$ respectively as described in Section 3. $\tau^{l}$ denotes the set of all nodes at level $l$ in tree $\mathcal{T}_{\chi}$ and $\nu^{l}$ denotes the set of all nodes at level $l$ in tree $\mathcal{T}_{\xi}$. Additionally, we assume that for each level $l\in\{0,\dots,L\}$, the ranks of all the blocks are equal, that is, $\forall i,j\ r^{l}_{i,j}=R^{l}$.
@@
A transformation to the index $j\rightarrow(j_{0},\dots,j_{L})$ is defined analogously. This leads to reshaping of the matrix $\bm{T}$ into an order $2d+2$ tensor, $\boldsymbol{\mathcal{T}}$, by representing each entry of the matrix $(i,j)$ by a $(2d+2)$- tuple $\underbrace{(i_{0},\dots,i_{d}}_{i},\underbrace{j_{0},\dots,j_{d})}_{j}$.
@@
It is worth mentioning that the reformulation of the butterfly decomposition described in Eq. 11 as a tensor network representation described in Eq. 19 is a valid tensor network only when the ranks of all the blocks at each level are the same, and the size of the leaf nodes is also the same. This is due to the fact that only in this case, we are able to stack the dense sub-blocks in butterfly factors and represent them as tensors as in Eq. 19. Also, this tensor network reformulation is not to be confused with approaches that leverage tensor low-rank compression algorithms (e.g., Tucker decomposition) to construct tensorized butterfly algorithms (see [42] for an example), which leads to a different tensor diagram compared with Fig. 2(b).
@@ 5 Tensor Completion with Butterfly Decomposition
Tensor completion problem is a well-studied problem from numerical optimization perspective with many existing algorithms such as alternating least squares (ALS) [71, 44], adaptive moment estimation (ADAM) optimization [43], and Riemannian optimization [46]. Next, we present the implementation and analyze the computational complexity of the proposed ALS and ADAM optimization method within butterfly tensor completion context in Section 5.1. Several practical considerations required to implement these algorithms are discussed in Section 5.2
@@ 5.1 Optimization for Tensor Completion of Butterfly Matrices
The proposed ALS and ADAM optimization methods for butterfly completion are described in this subsection. We focus on the explanation of the ALS algorithm and its complexity analysis, and briefly summarize the ADAM algorithm as many computational kernels are shared between them.
@@ 5.1.1 Alternating Least Squares (ALS)
ALS has proven to be a competitive method for tensor completion problems for various tensor decompositions [44]. The proposed algorithm proceeds by minimizing one factor at a time, summarized in Algorithm 1. For minimizing factors of the tensorized butterfly, the resulting least squares to solve for the first parameter tensor $\boldsymbol{\mathcal{S}}^{1}$ is given for each $({i_{0},\dots,i_{L}})$ and $(j_{0},\dots,j_{L})\in\Omega_{i_{0},\dots,i_{L}}$, as
$\sum_{r_{1}}s^{1}_{i_{0}\dots i_{L}r_{1}}a_{i_{0}\dots i_{L-1}j_{0}\dots j_{L}r}=t_{i_{0}\dots i_{L}j_{0}\dots j_{L}},$
(27) $a_{i_{0}\dots i_{L-1}j_{0}\dots j_{L}r}=\sum_{r_{2},\dots,r_{L+1}}(\prod_{m=1}^{L+1}s^{m+1}_{i_{0}\dots i_{L-m}j_{0}\dots j_{m-1}r_{m}r_{m+1}}),$
@@
All of the above least squares problems can be solved by solving for fibers of $\boldsymbol{\mathcal{S}}^{(1)}$ and $\boldsymbol{\mathcal{S}}^{L+2}$, and slices of $\boldsymbol{\mathcal{S}}^{l},l\in\{2,\dots,L+1\}$ in parallel, similar to tensor train completion [29]. We solve the least squares equations by using normal equations that result in small symmetric positive semi-definite systems of size $r\times r$ for $\boldsymbol{\mathcal{S}}^{1}$ and $\boldsymbol{\mathcal{S}}^{L+2}$ or $r^{2}\times r^{2}$ for $\boldsymbol{\mathcal{S}}^{l},\forall l\in\{2,\dots,L+1\}$, that can be solved in parallel efficiently [79, 81]. We summarize the normal equations method in Algorithm 1. This algorithm proceeds by iterating over each fiber of $\boldsymbol{\mathcal{S}}^{1}$ and $\boldsymbol{\mathcal{S}}^{L+2}$ and slice of $\boldsymbol{\mathcal{S}}^{l},l\in\{2,\dots,L+1\}$. This corresponds to lines 16, 6, and 26 and this loop can be parallelized as there is no dependency. These steps and their complexities are explained as follows.
Algorithm 1 ALS for Tensorized Butterfly Completion
@@
Comparing with (20), QTT can be viewed as a more aggressive compression format than butterfly, consisting of order $4$ tensor cores instead of butterfly’s order $L+3$ tensor cores. The diagrams of butterfly and QTT of the same tensor with $L=3$ are shown in 2(b) and (c). The QTT completion algorithm (based on extension of the tensor train completion in [29]) will be used as one of our reference algorithms in numerical experiments of Section 6 and is discussed in Algorithm 5. We also remark that Algorithm 1 reduces to low-rank matrix completion algorithm when $L=0$, which is used as one of our reference algorithms in numerical experiments, as well as the initial guess to the butterfly completion algorithms (see Section 5.2).
@@
5.1.2 ADAM Optimization
Adaptive moment estimation (ADAM) [43] is a first-order optimization algorithm that combines momentum and adaptive learning rates, making it effective for training deep models and non-convex optimization tasks. ADAM is used in tensor completion problems to efficiently update the factors by obtaining gradients via sparse tensor contractions. For the factors in tensorized butterfly, the gradients are computed by computing the sparse tensor contractions in s (22), (24), and (23), with $\boldsymbol{\mathcal{Z}}=\mathcal{P}_{\Omega}\big(\boldsymbol{\mathcal{X}}(\mathcal{S})-\boldsymbol{\mathcal{T}}\big)$. The cost of computing the gradients is $O(|\Omega|Lr^{2})$ and a detailed description of the sparse tensor contractions for gradients is summarized in Algorithm 3. Once gradients are computed, updates can be computed via the ADAM optimization method as summarized in Algorithm 2. ADAM uses scalar operations on parameters with the total cost of single iteration being $O(|\Omega|L^{2}r^{2}+Lnr^{2})$. Replacing $L=O(\log n)$ in the previous expression, the total cost of a single ADAM iteration is $O(|\Omega|r^{2}\log^{2}n+nr^{2}\log n)$.
5.2 Initial Guess Generation
Both the proposed ALS and ADAM optimization methods require an initial guess for the factors. Although random initial guesses are used for many existing tensor and matrix completion algorithms for data science and scientific computing applications, we remark that they do not work well for the proposed butterfly completion algorithms, particularly for applications involving highly oscillatory matrices. This is particularly because the larger number of parameters leading to a more complex optimization landscape. Additionally, as we will demonstrate in numerical experiments, we use $\mathcal{O}(|\Omega|)=\mathcal{O}(nr\log n)$ observed entries, making full recovery significantly more challenging and requiring an initial guess that is closer to the true solution in Frobenius norm than a random guess. We propose to first use the low-rank matrix completion algorithm (i.e., Algorithm 1 with $L=0$ and $r=R$, $R$ being a constant much smaller than the actual rank of the matrix). Neither $\mathcal{O}(|\Omega|)$ nor $r=R$ is sufficient for the low-rank completion algorithm to recover the underlying matrix with sufficient accuracy, that is, the relative error over observed entries defined in (30) stays over $0.5$. However, the resultant low-rank matrix provides a rough approximation of the target matrix and can serve as a better initial guess than a random guess for the butterfly completion algorithm. Our experimental observations suggest that obtaining a low rank decomposition with relative test error of about $0.9$ is enough to finally obtain a good butterfly matrix approximation. It is worth mentioning that the low-rank completion algorithm only requires $\mathcal{O}(t_{\max}|\Omega|R^{2}+nR^{3})$ operations.
Algorithm 3 ADAM for Tensor Completion for Butterfly Matrix
@@ 5.2 Initial Guess Generation
Once the initial low-rank matrix is computed, the low-rank matrix of rank $R$ needs to be converted to a butterfly matrix of $L$ levels and rank $r$, i.e., $\boldsymbol{\mathcal{S}}^{i},i=1,\ldots,L+2$, to use it as the initial guess. This low-rank to butterfly conversion can be performed efficiently using a randomized algorithm based on a randomized algorithm that uses matrix-vector products of the butterfly matrix [60]. In what follows, we briefly describe the randomized algorithm shown in Algorithm 4.The input to Algorithm 4 is the low rank matrix $\bm{X}=\bm{A}\bm{B}$ of rank $R$, desired numbers of butterfly levels $L$ and rank $r$, and an oversampling integer $p$. The algorithm sketches the relevant blocks of $\bm{X}$ using Kronecker-structured sketching matrices. Specifically, it employs Gaussian random matrices in Kronecker product with appropriately sized identity matrices. We express these operations in tensor form in Algorithm 4.
@@
The algorithm proceeds by first reshaping both the low rank factors to tensors using the reshape presented in (13). We then generate a random Gaussian tensor of appropriate size. The tensor is a reshaping of matrix of size $n\times(r+p)$ using (13). A tensor contraction with cost $\mathcal{O}(nrR)$ is computed in Line 5 to sketch the factors using the generated random Gaussian tensor. The column and row basis for outer-most levels can then be computed with QR factorization with column pivoting over the blocks of the resulting tensors denoted by QRCP in Line 8. Note that these blocks can be executed in parallel, and the QR factorization costs $\mathcal{O}(nr^{2})$ operations, and the blocks of factor tensors are updated to the upper triangular factors obtained from the QR factorization. Next, for each inner level $l=1,\ldots,H$, a new Gaussian random tensor of order L+2lL+2-l is generated. The tensors implicitly represent Gaussian random matrices obtained as the Kronecker product of identity matrices and appropriately sized Gaussian random matrices. The sketching and QR factorization are performed at Line 11, 14. These operations respectively cost $\mathcal{O}(nRr^{2})$, $\mathcal{O}(nr^{3})$ for each level $l$. Once all factor cores have been computed, $\boldsymbol{\mathcal{S}}^{H+1}$ is updated by contracting the resulting tensors in Line 18, which requires $\mathcal{O}(nRr\log n)$ operations. Note that this computation is equivalent to the computation done for the middle-level in Algorithm $4.1$ in [60] and sketching is not required for the middle level factor when the matrix is given as a low-rank decomposition. Overall, the complexity of Algorithm 4 is $\mathcal{O}(nRr\log n+nRr^{2}\log n+nr^{3}\log n)$.

### v3

@@ 3.3 Butterfly Decomposition in Machine Learning
Due to its ability to approximate Fourier integral operators with a small number of parameters, butterfly decomposition has recently been integrated into deep neural networks to efficiently learn structured linear transforms [dao2019learning, Fast-transform]. Building on this, the Kaleidoscope hierarchy, which is a product of learnable orthogonal butterfly matrices, was introduced into the Mel-Frequency Spectral Coefficients (MFSC) pipeline [davis1980comparison], achieving performance comparable to hand-crafted transforms. These Kaleidoscope matrices are capable of learning latent permutations and accelerating inference [dao2020kaleidoscope]. To further enhance their practicality and compatibility with GPU architectures, the sparsity structure of the butterfly factors was modified in various works [DEBUT, pmlr-v162-dao22a, chen2022pixelated]. Empirical evaluations in these works show that restructuring maintains similar accuracy while improving efficiency by enhancing the ease of implementation.
Integrating butterfly matrices into machine learning frameworks like PyTorch has led to new developments in neural network architectures and computational kernels. Although butterfly decompositions theoretically enable matrix-vector products in $O(n\log n)$ complexity, achieving these speedups in practice poses significant challenges. Traditional block-sparse representations often lead to inefficient high-level implementations, as they require manual looping over sparse blocks, which impedes performance. Realizing theoretical speedups for both the forward pass, that is, applying the butterfly matrix, and backward pass, which is obtaining the gradients of block sparse matrices, demands carefully optimized software kernels. While recent efforts have developed efficient GPU kernels for the forward pass [gonon2024make, fan2022adaptablebutterflyacceleratorattentionbased], kernels for computing gradients (i.e., the backward pass) remain underdeveloped. To address this limitation, some works have modified the sparsity structure of the butterfly factors to better align with hardware efficiency [pmlr-v162-dao22a]. In contrast, we propose a new representation of the butterfly decomposition as a tensor network, enabling differentiation with respect to the factors by leveraging automatic differentiation with tensor contractions. This reformulation is detailed in the next section.
@@ 4.1 Butterfly Decomposition Rewritten as a Tensor Network
To describe the tensor decomposition reformulation, we will continue to use the same assumptions as in the previous section. We consider an $m\times n$ input matrix $\bm{T}$, where $n=c\cdot 2^{L}$, and $L$ is the number of levels in the butterfly decomposition. The row and column sets, $\chi=\{0,\dots,m-1\}$ for rows and $\xi=\{0,\dots,n-1\}$ for columns of the matrix, and $m=n$. These sets are organized into hierarchical trees $\mathcal{T}_{\chi}$ and $\mathcal{T}_{\xi}$ respectively as described in Section 3. $\tau^{l}$ denotes the set of all nodes at level $l$ in tree $\mathcal{T}_{\chi}$ and $\nu^{l}$ denotes the set of all nodes at level $l$ in tree $\mathcal{T}_{\xi}$. Additionally, we assume that for each level $l\in\{0,\dots,L\}$, the ranks of all the blocks are equal, that is, $\forall i,j,\ r^{l}_{i,j}=R_{l}$. For the ease of later complexity analysis, we further assume $R_{l}=r$.
@@
A transformation to the index $j\rightarrow(j_{0},\dots,j_{L})$ is defined analogously. This leads to reshaping of the matrix $\bm{T}$ into an order $2L+2$ tensor, $\boldsymbol{\mathcal{T}}$, by representing each entry of the matrix $(i,j)$ by a $(2L+2)$- tuple $\underbrace{(i_{0},\dots,i_{d}}_{i},\underbrace{j_{0},\dots,j_{d})}_{j}$.
@@
It is worth mentioning that the reformulation of the butterfly decomposition described in Eq. 11 as a tensor network representation described in Eq. 19 is a valid tensor network only when the ranks of all the blocks at each level are the same, and the size of the leaf nodes is also the same. This is due to the fact that only in this case, we are able to stack the dense sub-blocks in butterfly factors and represent them as tensors as in Eq. 19. Also, this tensor network reformulation is not to be confused with approaches that leverage tensor low-rank compression algorithms (e.g., Tucker decomposition) to construct tensorized butterfly algorithms (see [kielstra2024linear] for an example), which leads to a different tensor diagram compared with Fig. 2(b).
@@ 5 Tensor Completion with Butterfly Decomposition
Tensor completion problem is a well-studied problem from numerical optimization perspective with many existing algorithms such as alternating least squares (ALS) [paatero1997223, kolda2009tensor], adaptive moment estimation (ADAM) optimization [kingma2014adam], alternating direction fitting (ADF) [grasedyck2015variants], and Riemannian optimization [kressner2014low]. Next, we present the implementation and analyze the computational complexity of the proposed butterfly completion algorithms using ALS, ADAM and ADF in Section 5.1. Several practical considerations required to implement these algorithms are discussed in Section 5.2
@@ 5.1 Optimization for Tensor Completion of Butterfly Matrices
The proposed ALS, ADAM and ADF optimization methods for butterfly completion are described in this subsection. We focus on the explanation of the ALS algorithm and its complexity analysis, and briefly summarize the ADAM and ADF algorithms as many computational kernels are shared between them.
@@ 5.1.1 Alternating Least Squares (ALS)
ALS has proven to be a competitive method for tensor completion problems for various tensor decompositions [kolda2009tensor]. The proposed ALS-based butterfly completion algorithm proceeds by minimizing one factor at a time, summarized in Algorithm 1. For minimizing factors of the tensorized butterfly, the resulting least squares to solve for the first parameter tensor $\boldsymbol{\mathcal{S}}^{1}$ is given for each $({i_{0},\dots,i_{L}})$ and $(j_{0},\dots,j_{L})\in\Omega_{i_{0},\dots,i_{L}}$, as
$\sum_{r_{1}}s^{1}_{i_{0}\dots i_{L}r_{1}}a_{i_{0}\dots i_{L-1}j_{0}\dots j_{L}r_{1}}=t_{i_{0}\dots i_{L}j_{0}\dots j_{L}},$
(27) $a_{i_{0}\dots i_{L-1}j_{0}\dots j_{L}r_{1}}=\sum_{r_{2},\dots,r_{L+1}}(\prod_{m=1}^{L+1}s^{m+1}_{i_{0}\dots i_{L-m}j_{0}\dots j_{m-1}r_{m}r_{m+1}}),$
@@
All of the above least squares problems can be solved by solving for fibers of $\boldsymbol{\mathcal{S}}^{(1)}$ and $\boldsymbol{\mathcal{S}}^{L+2}$, and slices of $\boldsymbol{\mathcal{S}}^{l},l\in\{2,\dots,L+1\}$ in parallel, similar to tensor train completion [grasedyck2015variants]. We solve the least squares equations by using normal equations that result in small symmetric positive semi-definite systems of size $r\times r$ for $\boldsymbol{\mathcal{S}}^{1}$ and $\boldsymbol{\mathcal{S}}^{L+2}$ or $r^{2}\times r^{2}$ for $\boldsymbol{\mathcal{S}}^{l},\forall l\in\{2,\dots,L+1\}$, that can be solved in parallel efficiently [singh2022distributed, smith2016exploration]. We summarize the normal equations method in Algorithm 1. This algorithm proceeds by iterating over each fiber of $\boldsymbol{\mathcal{S}}^{1}$ and $\boldsymbol{\mathcal{S}}^{L+2}$ and slice of $\boldsymbol{\mathcal{S}}^{l},l\in\{2,\dots,L+1\}$. This corresponds to lines 16, 6, and 26 and this loop can be parallelized as there is no dependency. These steps and their complexities are explained as follows.
Algorithm 1 ALS for Butterfly Completion
@@
Comparing with (20), QTT can be viewed as a more aggressive compression format than butterfly, consisting of order $4$ tensor cores instead of butterfly’s order $L+3$ tensor cores. The diagrams of butterfly and QTT of the same tensor with $L=3$ are shown in 2(b) and (c). The QTT completion algorithm (based on extension of the tensor train completion in [grasedyck2015variants]) will be used as one of our reference algorithms in numerical experiments of Section 6 and is discussed in Algorithm 5. We also remark that Algorithm 1 reduces to low-rank matrix completion algorithm when $L=0$, which is used as one of our reference algorithms in numerical experiments, as well as the initial guess to the butterfly completion algorithms (see Section 5.2).
5.1.2 ADAM Optimization
Adaptive moment estimation (ADAM) [kingma2014adam] is a first-order optimization algorithm that combines momentum and adaptive learning rates, making it effective for training deep models and non-convex optimization tasks. ADAM is used in tensor completion problems to efficiently update the factors by obtaining gradients via sparse tensor contractions. The proposed ADAM-based butterfly completion algorithm is described in Algorithm 3. For the factors in tensorized butterfly, the gradients are computed by computing the sparse tensor contractions in s (22), (24), and (23), with $\boldsymbol{\mathcal{Z}}=\mathcal{P}_{\Omega}\big(\boldsymbol{\mathcal{X}}(\mathcal{S})-\boldsymbol{\mathcal{T}}\big)$. The cost of computing the gradients is $O(|\Omega|Lr^{2})$ and a detailed description of the sparse tensor contractions for gradients is summarized in Algorithm 3. Once gradients are computed, updates can be computed via the ADAM optimization method as summarized in Algorithm 2. ADAM uses scalar operations on parameters with the total cost of single iteration being $O(|\Omega|L^{2}r^{2}+Lnr^{2})$. Replacing $L=O(\log n)$ in the previous expression, the total cost of a single ADAM iteration is $O(|\Omega|r^{2}\log^{2}n+nr^{2}\log n)$. Comparing with the total cost of a single ALS iteration, ADAM has an $\mathcal{O}(r^{2})$ reduction for the leading order term. That said, ADAM oftentimes requires a much larger number of iterations to converge compared with ALS. Just like ALS, we also implement a QTT completion algorithm with ADAM optimization with the same computational kernels as used for butterfly completion.
@@ 5.1.2 ADAM Optimization
Algorithm 3 ADAM for Butterfly Completion
@@
5.1.3 Alternating Direction Fitting (ADF)
Gradient-based methods like ADAM are computationally cheaper than ALS, however, these methods converge more slowly, taking much more time to reach an accurate decomposition. An alternative to solving least squares systems that arise in the ALS method described above is the ADF algorithm [grasedyck2013alternating]. ADF can be viewed as a variant of ALS that aims at reducing the cost per iteration while maintaining a similar convergence property to get the best of both worlds [grasedyck2015variants]. The ADF method relies on each of the linear systems being solved for ALS subiteration having orthonormal columns. Once this property is satisfied, ADF replaces these least-squares solves with gradient-based updates of individual core entries with accelerated convergence using an over-relaxation step size computation scheme. Our tensor network reformulation of the butterfly decomposition allows us to make the observation that this property can be satisfied for the butterfly decomposition, enabling a straightforward implementation of the ADF algorithm.
The step-size for each slice of the cores is determined by using the formula described in Algorithm $5$ in [grasedyck2015variants]. The key comparison point with respect to ALS is that ALS forms normal equations in Line 32 (and Lines 11 and 21) of Algorithm 1, leading to linear systems of sizes $r^{2}\times r^{2}$ requiring $\mathcal{O}(r^{4})$ cost for each slice. Instead, ADF does not compute these systems at all and works with gradient and sparse butterfly reconstruction contractions that cost $\mathcal{O}(|\Omega|r^{2}\log^{2}n)$ per iteration with an additional cost of orthogonalizing cores at each subiteration, which requires $\mathcal{O}(n)$ parallel QR decompositions of matrices of size $2r\times r$. The total per iteration cost of ADF for butterfly thus becomes $\mathcal{O}(|\Omega|r^{2}\log^{2}n+nr^{2}\log n)$. The implementation details of the orthogonalization for ADF for butterfly completion are described in Section A.3. The proposed ADF-based butterfly completion algorithm is described as Algorithm 6.
5.2 Initial Guess Generation
The ALS, ADAM and ADF methods require an initial guess for the factors. Although random initial guesses are used for many existing tensor and matrix completion algorithms for data science and scientific computing applications, we remark that they do not work well for the proposed butterfly completion algorithms ( and the reference QTT completion algorithms ), particularly for applications involving highly oscillatory matrices. This is particularly because the larger number of parameters leading to a more complex optimization landscape. Additionally, as we will demonstrate in numerical experiments, we use $|\Omega|=\mathcal{O}(nr\log n)$ observed entries, making full recovery significantly more challenging and requiring an initial guess that is closer to the true solution in Frobenius norm than a random guess. We propose to first use the low-rank matrix completion algorithm (i.e., Algorithm 1 with $L=0$ and $r=R$, $R$ being a constant much smaller than the actual rank of the matrix). Neither $|\Omega|=\mathcal{O}(nr\log n)$ nor $r=R$ is sufficient for the low-rank completion algorithm to recover the underlying matrix with sufficient accuracy, that is, the relative error over observed entries defined in (30) stays over $0.5$. However, the resultant low-rank matrix provides a rough approximation of the target matrix and can serve as a better initial guess than a random guess for the butterfly completion algorithm. Our experimental observations suggest that obtaining a low rank decomposition with relative test error of about $0.9$ is enough to finally obtain a good butterfly matrix approximation. It is worth mentioning that the low-rank completion algorithm only requires $\mathcal{O}(t_{\max}|\Omega|R^{2}+nR^{3})$ operations.
@@ 5.2 Initial Guess Generation
Once the initial low-rank matrix is computed, the low-rank matrix of rank $R$ needs to be converted to a butterfly matrix of $L$ levels and rank $r$, i.e., $\boldsymbol{\mathcal{S}}^{i},i=1,\ldots,L+2$, to use it as the initial guess. This low-rank to butterfly conversion can be performed efficiently using a randomized algorithm based on a randomized algorithm that uses matrix-vector products of the butterfly matrix [liu2021butterfly]. In what follows, we briefly describe the randomized algorithm shown in Algorithm 4.The input to Algorithm 4 is the low rank matrix $\bm{X}=\bm{A}\bm{B}$ of rank $R$, desired numbers of butterfly levels $L$ and rank $r$, and an oversampling integer $p$. The algorithm sketches the relevant blocks of $\bm{X}$ using Kronecker-structured sketching matrices. Specifically, it employs Gaussian random matrices in Kronecker product with appropriately sized identity matrices. We express these operations in tensor form in Algorithm 4.
The algorithm proceeds by first reshaping both the low rank factors to tensors using the reshape presented in (13). We then generate a random Gaussian tensor of appropriate size. The tensor is a reshaping of matrix of size $n\times(r+p)$ using (13). A tensor contraction with cost $\mathcal{O}(nrR)$ is computed in Line 5 to sketch the factors using the generated random Gaussian tensor. The column and row basis for outer-most levels can then be computed with QR factorization with column pivoting over the blocks of the resulting tensors denoted by QRCP in Line 8. Note that these blocks can be executed in parallel, and the QR factorization costs $\mathcal{O}(nr^{2})$ operations, and the blocks of factor tensors are updated to the upper triangular factors obtained from the QR factorization. Next, for each inner level $l=1,\ldots,H$, a new Gaussian random tensor of order L+2lL+2-l is generated. The tensors implicitly represent Gaussian random matrices obtained as the Kronecker product of identity matrices and appropriately sized Gaussian random matrices. The sketching and QR factorization are performed at Line 11, 14. These operations respectively cost $\mathcal{O}(nRr^{2})$, $\mathcal{O}(nr^{3})$ for each level $l$. Once all factor cores have been computed, $\boldsymbol{\mathcal{S}}^{H+1}$ is updated by contracting the resulting tensors in Line 18, which requires $\mathcal{O}(nRr\log n)$ operations. Note that this computation is equivalent to the computation done for the middle-level in Algorithm $4.1$ in [liu2021butterfly] and sketching is not required for the middle level factor when the matrix is given as a low-rank decomposition. Overall, the complexity of Algorithm 4 is $\mathcal{O}(nRr\log n+nRr^{2}\log n+nr^{3}\log n)$.